Write a quadratic equation in the variable x having the given numbers as solutions. Type the equation in standard form, ax^2 + bx + c = 0.
- sqrt( 3 ) root( 3, 3 )

Answer by TakeATuition.com(57)

solution -
The form of quadratic equation is
ax^2 + bx + c = 0
if two roots of the equation are given then
equation can be formed by this formula
x^2 - (sum of the roots)x + product of the roots = 0 .....(i)

Now two roots are given that is 3 and 3
Therefore
sum of the roots = 3 + 3 = 6
product of the roots = 3*3 = 9
substituting the value in equation (i)
we have
x^2 - (6)x + 9 = 0
x^2 - 6x + 9 = 0
is the required equation.
FOR MORE HELP VISIT : www.TakeATuition.com